OpenStudy (anonymous):

Verify the equation secx-1/1-cosx=secx

OpenStudy (anonymous):

is this \[(\sec(x)-1)/(1-\cos(x))=\sec(x)\] ?

OpenStudy (anonymous):

yes how do you verify that?

OpenStudy (anonymous):

try and use identities to prove that the first side equals the 2nd, i am trying to figure it out.

OpenStudy (anonymous):

just put sec x = 1/ cos x then (1-cos x) will cancel out from numerator and denominator then only 1/cos x will remain. which is sec x

OpenStudy (anonymous):

show me how you did that

OpenStudy (anonymous):